# StructuredField: Unifying Structured Geometry and Radiance Field

### v1

@@ 3.2 Inversion-free Structured Geometry
Figure 6: Invertible neural network architecture. By applying a permutation strategy, the input vector is divided into two contiguous parts, ensuring that the resulting Jacobian matrix is upper triangular with positive diagonal elements. This strategy guarantees that the determinant of the Jacobian is positive, thus ensuring that the mapping is inversion-free.
@@
  • Homeomorphism guarantees that tetrahedral mesh after optimization does not produce self-intersecting boundaries.
During optimization, both the topology $\mathcal{T}$ and vertex positions $\mathcal{V}$ of tetrahedral mesh are fixed. The orientation-preserving homeomorphism maps a vertex $v=[x,y,z]$ to its actual position as follows:
$v^{\prime}=[x^{\prime},y^{\prime},z^{\prime}]^{T}=\mathcal{H}([x,y,z]^{T}).$ (6)
@@
Based on these observations, we implement the orientation-preserving homeomorphism $\mathcal{H}$ by a novel invertible network. Real-NVP [12] appears to be a reasonable network architecture, as it is bijective and ensures that the Jacobian determinant of each coupling layer remains positive. Real-NVP splits the input $x\in\mathbb{R}^{D}$ into two components, $x_{1:d}$ and $x_{d+1:D}$. The output of a coupling layer $x^{\prime}\in\mathbb{R}^{D}$ follows the equations:
$x^{\prime}_{1:d}$ $=x_{1:d}$ (7)
$x^{\prime}_{d+1:D}=x_{d+1:D}$ $\cdot\exp(s(x_{1:d}))+t(x_{1:d}),$
@@
$J_{h}=\begin{bmatrix}1&0\\ \frac{\partial x^{\prime}_{d+1:D}}{\partial x^{T}_{1:d}}&\text{diag}[\exp(s(x_{1:d})]\end{bmatrix}.$ (8)
@@
$J_{h}=\begin{bmatrix}1&\frac{\partial y^{\prime}}{\partial x}&0\\ 0&\exp(s)&0\\ 0&\frac{\partial y^{\prime}}{\partial z}&1\end{bmatrix}$ (9)
@@
$\phi\left([x,y,z]^{T}\right)$ $=P^{-1}\left(h\left(P[x,y,z]^{T}\right)\right)$ (10)
$P=$ $\begin{bmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{bmatrix}.$
@@ 3.3 Training
We use two different initialization strategies. In the first strategy, we reconstruct the surface using NeuS2 [55], and then generate the tetrahedral mesh from the reconstructed triangle mesh using fTetWild [21]. The second is to generate a uniform tetrahedral grid within the bounding box of the scene. We conduct experiments on different initialization strategies in Sec. 4.1, which show that even with poorly initialized tetrahedral mesh, we can still achieve high-quality results.
Adaptive Tetrahedron Control
We design an adaptive control strategy to dynamically add and remove tetrahedra. Specifically, if the gradient of a tetrahedron’s corresponding Gaussians exceeds a certain threshold $\delta$, the tetrahdron is split. We add new vertices at the center of the tetrahedron and at the midpoints of each edge, thereby subdividing the tetrahedron into eight smaller tetrahedra. On the other hand, tetrahedra with corresponding Gaussians’ opacities less than $\epsilon$ are pruned.
@@
Our model is trained using a loss function written as a weighted sum:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\mathrm{1}}+\lambda_{2}\mathcal{L}_{\mathrm{SS}\mathrm{IM}}+\lambda_{3}\mathcal{L}_{\text{mask}}.$ (11)
The loss used for supervision of the RGB signal: $\mathcal{L}_{\mathrm{1}}$ and $\mathcal{L}_{\mathrm{SSIM}}$ follow 3D Gaussian splatting [30]. $\mathcal{L}_{mask}$ is designed to prevent artifacts at the object boundaries:
@@
4 Eexperiments
We first give the implementation details. The contribution analysis of proposed components is given in Sec. 4.1. Evaluations of our proposed representation, with previous state-of-the-art non-editable and editable approaches are presented in Sec. 4.2. We also show some applications of our proposed representation in Sec. 4.3
@@ 4 Eexperiments
We set $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.5$ for training the model. We use the ReLU function to prevent the weights of the vertices from being negative, and apply the sigmoid function to ensure that the opacity stays within the range of $[0,1)$. The gradient threshold $\delta$ for tetrahedron splitting is set to 0.0002, the same value used in 3DGS. We set the pruning threshold $\epsilon=0.0005$ to prevent overly aggressive pruning that could result in too many isolated tetrahedra. The invertible neural network consists of 3 blocks, with each axis chosen in order. A 2D hash encoding is used with a size of $2^{19}$ and a maximum resolution of $1024^{2}$. The MLP has 2 hidden layers, each with 128 units. Additionally, we reconstruct the surface with NeuS2 [55] and generate the initial tetrahedral mesh through fTetWild [21]. We train each model with 30,000 iterations, which takes approximately 1 hour. The final model has an average of 826k tetrahedra with 197k vertices (for the NeRF Synthetic dataset). All experiments are conducted on a single GTX RTX4090 GPU.
4.1 Abalation Study
@@
  • Q1: How much does our orientation-preserving homeomorphism affect rendering quality and mesh quality?
@@
Table 1: Quantitative results with different constraints.
PSNR ↑ SSIM ↑ LPIPS ↓
w/o constraint 30.41 0.947 0.062
w/ $\mathcal{L}_{volume}$ 29.37 0.925 0.076
w/ $\mathcal{L}_{sv}$ 29.92 0.936 0.069
ours 30.23 0.944 0.065
Q1: Effect of Orientation-preserving Homeomorphism.
We conduct this experiment using the NeRF synthetic dataset, with the mesh initialized by NeuS2 as input. To avoid the influence of adaptive control on the experiment, we removed this module, fixing the number of tetrahedra. We regularized the training of the tetrahedral mesh using four different methods: without constraints, with $\mathcal{L}_{\text{volume}}$, with $\mathcal{L}_{\text{sv}}$, and with our orientation-preserving homeomorphism as a constraint. Tab. 1 shows the quantative results for the reconstruction of these four methods. Compared to the unconstrained optimization, our module has minimal impact on the rendering quality, whereas the other regularization terms significantly degrade the reconstruction quality. As shown in Fig. 7, under unconstrained conditions, the mesh exhibits a large number of overlaps and abnormal structures. Additionally, $\mathcal{L}_{\text{sv}}$ and $\mathcal{L}_{\text{volume}}$ constraints fail to effectively regulate the mesh structure. Since the local orientation-preserving property near the vertices cannot guarantee global inversion-free tetrahedra, our final result still contains a small number of inverted elements. Nonetheless, our method produces much better tetrahedral meshe compared to the constraints imposed by the loss function and achieves similar reconstruction quality as the unconstrained approach.
Table 2: Quantitative results for different initializations.
@@
w/ uniform 33.19 0.969 36.12 0.975 35.78 0.986
w/ NeuS2 35.48 0.982 37.69 0.985 37.56 0.993
Figure 7: Visualization of tetrahedral mesh optimized with different constraints.
@@
We conduct this experiment on three scenes from the NeRF synthetic dataset: Lego, Hotdog, Mic. Initially, we start with a uniform tetrahedral mesh as described in Sec. 3.3, and project the tetrahedra onto the image plane, followed by tetrahedron pruning according to the mask. This tetrahedral mesh is then used as input to our method for training. Tab. 2 shows the final reconstruction results, compared to those obtained with NeuS2 initialization. Our method still produces good mesh quality and comparable reconstruction results, indicating that our approach is robust to mesh initialization.
4.2 Comparison
Figure 8: Our results on the test-views of Shelly dataset and NeRF Synthetic dataset.
We conduct experiments to compare the performance of rendering on static scenes. The results show that, despite the need to maintain a structured geometry during optimization, we are still able to achieve comparable result to that of unstructured radiance field representation like 3DGS [30].
Table 3: Quantitative results on NeRF Synthetic and Shelly dataset.
@@ 4.2 Comparison
3DGS 33.78 0.969 0.030 39.61 0.961 0.064
2DGS 32.92 0.966 0.036 34.62 0.928 0.098
Mani-GS 32.58 0.961 0.034
SuGaR 29.78 0.948 0.055 31.19 0.943 0.080
Ours 33.53 0.968 0.019 39.76 0.967 0.052
We employ two datasets for evaluation: NeRF Synthetic dataset [39], containing eight scenes, and Shelly dataset [57], containing six scenes. We conduct comparisons with implicit, point-based, and mesh-based methods to evaluate the performance of our approach, including NeRF [39], 3DGS [30], 2DGS [22], Adaptive Shells [57] and Mani-GS [17]. We use surface reconstructed by NeuS2 as input to Mani-GS for a fair comparison. We use the same training process as 3DGS. The rendering quality is evaluated using a set of established metrics: PSNR, SSIM [56], and LPIPS [66]. In Tab. 3, we conduct a quantitative comparison of rendering quality. Our approach shows competitive results across all mesh-based and point-based methods according to three evaluation metrics. Moreover, our approach achieves qualitatively better reconstructions, with fewer artifacts and more detailed results, as shown in Fig. 8.
4.3 Applications
Figure 9: Gallery of physical simulation results. The original shape is shown in the leftmost column, with the corresponding physical simulations displayed in the right column. Videos are provided in the supplementary materials for further visualization.
Figure 10: Gallery of deformation results. Videos are provided in the supplementary materials for further visualization.
Our method directly constructs an explicit tetrahedral mesh $\mathcal{M}$ with good quality, and achieves good rendering quality at the same time, which proves highly valuable for a variety of downstream applications. With tetrahedral mesh, we can perform physics simulations, animations, editing, and other operations seamlessly. With our proposed StructuredField, we can easily update the mesh by simply tracking the movement or displacement of each vertex. This allows us to reparameterize the Gaussians based on the new positions of vertices, enabling the rendering of deformed scenes.

### v2

@@ 3.2. Hierarchical Implicit subdivision
Figure 4. Invertible neural network architecture. By applying a permutation strategy, the input vector is divided into two contiguous parts, ensuring that the resulting Jacobian matrix is upper triangular with positive diagonal elements. This strategy guarantees that the determinant of the Jacobian is positive, thus ensuring that the mapping is inversion-free.
High-fidelity rendering often requires capturing details at multiple scales, a capability demonstrated by many successful 3D reconstruction methods. For example, Instant-NGP (Müller et al., 2022) uses multi-resolution hash encodings, combining features from various grid resolutions to efficiently represent intricate, multi-scale details. Similarly, 3D Gaussian Splatting (Kerbl et al., 2023) employs adaptive density control, refining Gaussian primitives in under reconstructed regions to better capture fine-scale information. Therefore, to equip our tetrahedral representation with a comparable capability to capture details across diverse scales, a mechanism for refining the tetrahedral elements themselves is essential.
While explicitly subdividing tetrahedra in the base mesh $\mathcal{M}$ can introduce finer geometric details, this direct modification of the mesh topology often creates non-conformal interfaces or degenerate elements. Such non-conformality and the presence of degenerate or ill-conditioned elements are highly problematic for downstream applications like physical simulation, which demand well-structured, conformal meshes, and also invariably increase the complexity of the simulation mesh. To circumvent these drawbacks and equip our representation with the capability to capture details at different scales, we propose a hierarchical implicit subdivision strategy.
Hierarchical Implicit Subdivision
During optimization, base tetrahedra in $\mathcal{M}$ identified for refinement undergo 1-to-4 subdivision, initiating the formation of a tree hierarchy. For each subdivided tetrahedron, the split is achieved by introducing a new, optimizable control point conceptually located within it. The position of this control point is parameterized by its barycentric coordinates relative to its respective parent tetrahedron. These coordinates are optimized during training, ensuring the control point remains within the parent’s volume while allowing for adaptive shaping of the resulting child tetrahedra. This subdivision process can be applied recursively if multiple levels of detail are desired, further extending this hierarchical structure.
This hierarchical structure fulfills a crucial dual role in our representation. The root nodes of this tree consistently form the conformal mesh $\mathcal{M}$ designated for physical simulation, preserving its fixed topology and structural integrity. Conversely, the leaf tetrahedra of this tree hierarchy, representing the finest level of adaptive detail, constitute the mesh $\mathcal{M}_{r}$ utilized for high-fidelity rendering.
3.3. Inversion-free Structured Geometry
One of our goals is to optimize the base mesh $\mathcal{M}$ while maintaining high-quality structure. As shown in Fig. 6 (W/o constraints), optimizing a tetrahedral mesh without any constraints leads to poor mesh quality. The main cause of this issue is that the vertices can freely move during the optimization of the tetrahedral mesh. Consequently, the tetrahedral mesh suffers from self-intersections and element inversion problems, resulting in tetrahedra overlaps as shown in Fig. 5.
Figure 5. Two causes leading to mesh anomalies: self intersection and element inversion.
A naive solution is to introduce a loss function to constrain the mesh optimization process. Locally injective mapping (Schüller et al., 2013) uses a barrier function to strictly prevent element inversion. However, the gradient explosion near the critical points of the barrier function makes it difficult to use directly in gradient descent-based optimization. Another approach is to apply an $L_{1}$-norm constraint on the signed volume of the optimized tetrahedra $L_{\text{sv}}$, making tetrahedra with negative signed volume sparse:
(5) $\mathcal{L}_{\text{sv}}$ $=\frac{1}{K}\sum_{k=1}^{K}\max\left(-s(T_{k}),0\right),$
where $s(T_{k})$ is the signed volume of the tetrahedron $T_{k}$, $T^{0}_{k}$ is the initial tetrahedron. However, we found that these loss function-based constraints significantly impact the rendering quality and do not fully guarantee that the tetrahedral mesh remains a reasonable structure as shown in Sec. 4.1.
@@ 3.3. Inversion-free Structured Geometry
  • Homeomorphism guarantees that the tetrahedral mesh after optimization does not produce self-intersecting boundaries.
During optimization, both the topology $\mathcal{T}$ and vertex positions $\mathcal{V}$ of the base mesh $\mathcal{M}$ are fixed. The orientation-preserving homeomorphism maps a vertex $v=[x,y,z]$ to its actual position as follows:
(6) $v^{\prime}=[x^{\prime},y^{\prime},z^{\prime}]^{T}=\mathcal{H}([x,y,z]^{T}).$
@@
Based on these observations, we implement the orientation-preserving homeomorphism $\mathcal{H}$ by a novel invertible network. Real-NVP (Dinh et al., 2017) appears to be a reasonable network architecture, as it is bijective and ensures that the Jacobian determinant of each coupling layer remains positive. Real-NVP splits the input $x\in\mathbb{R}^{D}$ into two components, $x_{1:d}$ and $x_{d+1:D}$. The output of a coupling layer $x^{\prime}\in\mathbb{R}^{D}$ follows the equations:
(7) $x^{\prime}_{1:d}=x_{1:d},\quad x^{\prime}_{d+1:D}=x_{d+1:D}\cdot\exp(s(x_{1:d}))+t(x_{1:d}),$
@@
(8) $J_{h}=\begin{bmatrix}I&0\\ \frac{\partial x^{\prime}_{d+1:D}}{\partial x^{T}_{1:d}}&\text{diag}[\exp(s(x_{1:d})]\end{bmatrix}$.
@@
(9) $J_{h}=\begin{bmatrix}1&\frac{\partial y^{\prime}}{\partial x}&0\\ 0&\exp(s)&0\\ 0&\frac{\partial y^{\prime}}{\partial z}&1\end{bmatrix}$
@@
(10) $\phi\left([x,y,z]^{T}\right)=P^{-1}\left(h\left(P[x,y,z]^{T}\right)\right),\quad P=\begin{bmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{bmatrix}$.
@@ 3.4. Training
We use two different initialization strategies. In the first strategy, we reconstruct the surface using NeuS2 (Wang et al., 2023a), and then generate the tetrahedral mesh from the reconstructed triangle mesh using fTetWild (Hu et al., 2020). The second is to generate a uniform tetrahedral grid within the bounding box of the scene. We conduct experiments on different initialization strategies in Sec. 4.1, which show that even with poorly initialized tetrahedral mesh, we can still achieve high-quality rendering results.
Adaptive Control of Primitives
As detailed in Sec. 3.2, our primary mechanism for adapting geometric detail for rendering is the hierarchical implicit subdivision strategy. This approach inherently avoids direct modification of the base mesh $\mathcal{M}$ topology. To manage the set of active rendering primitives in $\mathcal{M}_{r}$ during optimization, we introduce an adaptive masking mechanism. Specifically, rendering tetrahedra whose opacity $o_{k}$ falls below a predefined threshold $\epsilon$ are masked. These masked tetrahedra no longer contribute to the rendering process. This strategy contrasts with directly deleting tetrahedra from the fixed-topology base mesh, which could introduce problematic structural changes or compromise conformality.
@@
Our model is trained by minimizing a composite loss function:
(11) $\mathcal{L}=\mathcal{L}_{render}+\lambda_{3}\mathcal{L}_{\text{mask}}+\lambda_{4}\mathcal{L}_{\text{quality}}.$
The loss used for supervision of the RGB signal: $\mathcal{L}_{render}=\lambda_{1}\mathcal{L}_{\mathrm{1}}+\lambda_{2}\mathcal{L}_{\mathrm{SSIM}}$ follow 3D Gaussian splatting (Kerbl et al., 2023). $\mathcal{L}_{mask}$ is designed to prevent artifacts at the object boundaries:
@@
To specifically enhance the geometric quality of the base tetrahedral mesh, we introduce $\mathcal{L}_{\text{quality}}$ (Lo, 1997):
(12) $\mathcal{L}_{\text{quality}}=\frac{1}{K}\sum_{k=1}^{K}max(r-Q,0),\quad Q=\frac{C_{0}V}{S_{\textup{rms}}},$
where $V$ represents the tetrahedron’s volume, $C_{0}$ is a normalization constant to ensure $Q=1$ for an equilateral tetrahedron, and $S_{rms}$ denotes the root-mean-square of the six edge lengths of the tetrahedron. This loss essentially constrains the quality of the tetrahedron to be greater than $r$.
4. Experiments
We first give the implementation details. The contribution analysis of proposed components is given in Sec. 4.1. Evaluations of our proposed representation with previous state-of-the-art non-editable and editable approaches are presented in Sec. 4.2. We also show some applications of our proposed representation in Sec. 4.3.
@@ 4. Experiments
We set $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, $\lambda_{3}=0.5$, and $\lambda_{4}=10.0$ for training the model. We use the ReLU function to prevent the weights of the vertices from being negative, and apply the sigmoid function to ensure that the opacity stays within the range of $[0,1)$. The gradient threshold $\delta$ for tetrahedron splitting is set to 0.0002, the same value used in 3DGS. We set the pruning threshold $\epsilon=0.05$ to mask low-opacity tetrahedra. The invertible neural network consists of 3 blocks, with each axis chosen in order. A 2D hash encoding is used with a size of $2^{19}$ and a maximum resolution of $1024^{2}$. The MLP has 2 hidden layers, each with 128 units. Additionally, we reconstruct the surface with NeuS2 (Wang et al., 2023a) and generate the initial tetrahedral mesh through fTetWild (Hu et al., 2020). Furthermore, we limit the maximum subdivision depth for each tetrahedron to 5 levels. We train each model with 30,000 iterations, which takes approximately 1 hour. All experiments are conducted on a single NVIDIA RTX4090 GPU.
4.1. Ablation Study
@@
  • Q1: How much does our orientation-preserving homeomorphism and $\mathcal{L}_{\text{quality}}$ affect rendering quality and mesh quality?
@@
Table 1. Quantitative results with different constraints. Mesh quality is measured by the Aspect Ratio Gamma.
Rendering Quality Mesh Quality
PSNR ↑ SSIM ↑ LPIPS ↓ Quality ↑ # Inverted ↓
(a) W/o Constraints 34.89 0.978 0.0210 $0.159\pm 0.179$ $\sim 83.2k$
(b) W/ $\mathcal{L}_{sv}$ 33.59 0.975 0.0294 $0.144\pm 0.187$ $\sim 53.6k$
(c) W/ $\mathcal{H}$ 34.64 0.974 0.0220 $0.471\pm 0.259$ $\sim 1.3k$
(d) W/ $\mathcal{H}$ + $\mathcal{L}_{quality}$ 33.64 0.974 0.0455 $0.816\pm 0.058$ 0
(e) W/o Optimization 32.78 0.971 0.0545 $0.797\pm 0.111$ 0
Figure 6. Visualization of tetrahedral mesh cross-sections optimized with different constraints.
Q1: Effect of Orientation-preserving Homeomorphism and $\mathcal{L}_{\textrm{quality}}$.
We conduct this experiment on the Lego and Chair scenes of NeRF synthetic dataset, with the mesh initialized by NeuS2 as input. To avoid the influence of number of primitives on the experiment, we remove the module of adaptive control of primitives, fixing the number of tetrahedra. We regularize the training of the tetrahedral mesh using four different methods: (a) without explicit constraints, (b) with the signed volume loss $\mathcal{L}_{\text{sv}}$, (c) with our orientation-preserving homeomorphism $\mathcal{H}$ as a constraint, and (d) with both $\mathcal{H}$ and our mesh quality loss $\mathcal{L}_{\text{quality}}$ applied together. Furthermore, we provide a (e) baseline where the tetrahedral mesh vertices are not optimized, and only other rendering attributes are refined.
Tab. 1 and Fig. 6 present a comparison of reconstruction quality and geometric quality for these methods. Optimizing the mesh directly (without constraints) and optimizing with our homeomorphism $\mathcal{H}$ achieve comparable rendering quality. However, employing our $\mathcal{H}$ results in significantly fewer inverted elements and higher overall geometric quality. In contrast, applying the $\mathcal{L}_{\text{sv}}$ loss has a limited effect on reducing the number of inverted elements and noticeably degrades rendering quality. This suggests that our approach of defining the feasible optimization space for vertices via an orientation-preserving homeomorphism is less susceptible to converging to poor local minima compared to methods based solely on regularization penalties. The geometric quality regularization term $\mathcal{L}_{\text{quality}}$ leads to a slight decrease in rendering quality compared to using $\mathcal{H}$ alone, but it yields a substantial improvement in the geometric quality of the mesh elements.
We present more visual results of the optimized mesh in Fig. 12 and numerical results in Tab. 5 and Tab. 5. The reported mesh quality metrics include AR (Aspect Ratio), defined as $3r_{k}/\rho_{k}$ (where $r_{k}$ is the inradius, and $\rho_{k}$ is the circumradius of tetrahedron $T_{k}$). ARG (Aspect Ratio Gamma) is consistent with the quality metric $Q$ in Eq. 12. Lastly, AVR (Adjacent cells Volume Ratio) measures the ratio of the largest to smallest cell volume within the 1-ring neighborhood of a tetrahedron. Notably, after constrain with our orientation-preserving homeomorphism $\mathcal{H}$ and $\mathcal{L}_{\text{quality}}$ term, the geometric quality of our final mesh often surpasses that of the initial mesh.
Table 2. Quantitative results for different initializations.
@@
W/o Rotation 33.27 0.968 36.14 0.975 36.06 0.986
W/ Uniform 34.98 0.980 36.90 0.983 37.46 0.991
W/ NeuS2 35.50 0.983 37.80 0.986 37.64 0.993
@@
We conduct this experiment on the NeRF synthetic dataset: lego, hotdog, mic. Initially, we start with a uniform tetrahedral mesh as described in Sec. 3.4. This tetrahedral mesh is then used as input into our method for training. Tab. 2 shows the final reconstruction results under different settings compared to those obtained with NeuS2 initialization. Notably, “W/o Rotation” means the removal of the optimizable rotation parameters in Eq. 3, which leads to a significant degradation in reconstruction quality. This performance drop is likely because the per-vertex weights are then solely responsible for determining both the shape and effective orientation of the 3D Gaussians, thereby limiting their expressive ability. Conversely, when the optimizable rotation parameters are included, the “W/ Uniform” initialization setting still produces reconstruction results comparable to those achieved with the NeuS2 initialization (“W/ NeuS2 ”).
4.2. Comparison
We conduct experiments to compare the performance of rendering on static scenes. The results show that, despite the need to maintain a structured geometry during optimization, we are still able to achieve comparable results to that of unstructured radiance field representation like 3DGS (Kerbl et al., 2023).
Table 3. Quantitative results on NeRF Synthetic and Shelly dataset.
@@ 4.2. Comparison
[…]
Table 4. Quantitative results on Mip-NeRF360 dataset. Results of other methods are borrowed from Radiant Foam’s paper.
Outdoor Scene Indoor Scene
PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓
[…]
Figure 7. Gallery of Rendering Results on MipNeRF 360 dataset.
We employ two datasets for evaluation: NeRF Synthetic dataset (Mildenhall et al., 2020a), containing eight scenes and Shelly dataset (Wang et al., 2023b), containing six scenes. We conduct comparisons with implicit, point-based, and mesh-based methods to evaluate the performance of our approach, including NeRF (Mildenhall et al., 2020a), 3DGS (Kerbl et al., 2023), 2DGS (Huang et al., 2024b), Adaptive Shells (Wang et al., 2023b) and Mani-GS (Gao et al., 2024). We use the surface reconstructed by NeuS2 as input to Mani-GS for a fair comparison. We use the same training process as 3DGS. The rendering quality is evaluated using a set of established metrics: PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018). In Tab. 3, we conduct a quantitative comparison of rendering quality. Our approach shows competitive results across all mesh-based and point-based methods according to three evaluation metrics. Moreover, our approach achieves qualitatively better reconstructions, with fewer artifacts and more detailed results, as shown in Fig. 8.
Furthermore, to demonstrate the effectiveness of our method in real-world unbounded scenes, we conducted experiments on the MipNeRF 360 dataset (Barron et al., 2022). For initialization in these scenes, we utilize a mesh of the centrally located region, initially reconstructed by Gaussian Opacity Fields (Yu et al., 2024). This mesh is then tetrahedralized using fTetWild (Hu et al., 2020) to serve as the input tetrahedral mesh for our method. Our StructuredField representation is employed to reconstruct this central region of interest, while the surrounding background is modeled using standard 3D Gaussian Splatting  (Kerbl et al., 2023). Tab. 4 presents a numerical comparison of our method against several state-of-the-art approaches, including 3DGS (Kerbl et al., 2023), Mip-Splatting (Yu et al., 2004), MipNeRF360 (Barron et al., 2022), and RadiantFoam (Govindarajan et al., 2025). Our method achieves reconstruction results comparable to those of 3DGS and surpasses RadiantFoam, another approach that also utilizes a deformable tetrahedral grid as its geometric primitive for scene representation. Additional visual results from our method on these scenes are provided in Fig. 7.
4.3. Applications
Figure 8. Our results on the test-views of Shelly dataset and NeRF Synthetic dataset.
Figure 9. Gallery of deformation results. Videos are provided in the supplementary materials for further visualization.
Our method directly constructs an explicit tetrahedral mesh $\mathcal{M}$ with good quality and achieves good rendering quality at the same time, which proves highly valuable for a variety of downstream applications. With tetrahedral mesh, we can perform physics simulations, animations, editing, and other operations seamlessly. With our proposed StructuredField, we can easily update the mesh by simply tracking the movement or displacement of each vertex. This allows us to reparameterize the Gaussians based on the new positions of vertices, enabling the rendering of deformed scenes.